Number of solutions of pair x, y of the equation sin x. sin y = min {–1, a2 - 4a + 5}, a belongs to R where 0 π, -π 

Grade:11

1 Answers

Anish Singhal
5 years ago
Out put of rhs will always be -1, draw graph of both. Now equate sinx.siny=-1 in interval -pi to pi, so (x=pi/2 and y=pi/2 )or (x=-pi/2 and y=pi/2)